OpenStudy (anonymous):

A surveying crew is tasked to measure the height of a mountain. From a point on level ground, they measure the angle of elevation to the top of the mountain as 21(degrees)40(minutes). They move 500m closer and found that the angle of elevation is now 35(degrees)10(minutes). How high is the mountain? I don't really need an answer, I just want to ask if this question involves both oblique triangles and right triangles to solve. Anyone care to verify this?

OpenStudy (anonymous):

tan(21d40m)=h/(500+x), tan35=h/x, h=h (500+x)tan(21d40m)=xtan35 500tan(21d40m)=xtan35-xtan(21d40m) 500tan(21d40m)=x[tan35-tan(21d40m)] 500tan(21d40m)/[tan35-tan(21d40m)]=x x=655.71 meters

OpenStudy (anonymous):

Entropy, let me know if you have question

OpenStudy (anonymous):

That's clever bud, you equated the both h since the trig ratios for them would produce the same height right?

OpenStudy (anonymous):

yeah, i think you can also use cosine law or tangent law

OpenStudy (anonymous):

I thought it should be sine law

OpenStudy (anonymous):

thanks and good luck now Entropy

OpenStudy (anonymous):

You got two angles and a side (500m).

OpenStudy (anonymous):

yes its sine law im sorry..lol

OpenStudy (anonymous):

No problem sir, thanks anyway.